Please help me with this! Determine the quadratic equation for the parabola:
Minimum value is 4 when x=3
zeros of f: one is 6

Answer by venugopalramana(3286)

Please help me with this! Determine the quadratic equation for the parabola:
Minimum value is 4 when x=3
zeros of f: one is 6
STD.EQN.IS
Y-K=A(X-H)^2
MINIMUM IS AT X=3...SO...H=3....AND HENCE K=4
Y-4=A(X-3)^2.....Y=0...WHEN...X=6...SO....
0=4+A(6-3)^2=4+9A
A=-4/9
Y-4=-4(X-3)^2/9
9Y-36+4(X-3)^2=0